OpenStudy (anonymous):

Simplify root(4) (w^4). Assume that w>0. FOR MEDALLLL

OpenStudy (anonymous):

is the w in the root?

OpenStudy (anonymous):

yes @PeachRings

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

well, do you know the square root of 4? :)

OpenStudy (anonymous):

2 :)

jigglypuff314 (jigglypuff314):

Correct! :D then do you know what the square root of w^4 is? for example square root of w^2 = w square root of w^6 = w^3 square root of w^8 = w^4 etc

OpenStudy (anonymous):

w^2 ?

jigglypuff314 (jigglypuff314):

Yes! :)

OpenStudy (anonymous):

so that's the answer than right ?

jigglypuff314 (jigglypuff314):

yes :) combine the two parts